An alloy is composed of 40 kg zinc and 10 kg of copper; another alloy contains 80 kg copper and 20 kg of tin. If three parts by mass of the first alloy are fused with one part by mass of the second alloy to make a third alloy, what percentage of copper and tin will be found in the third alloy?
question got from BTEC First MATHEMATICS for Technicians

Answer by htmentor(1343)

1st alloy contains 40/(40+10) = 4/5 zinc and 10/(40+10) = 1/5 copper
2nd alloy contains 80(80+20) = 4/5 copper and 20(80+20) = 1/5 tin
If we make a mixture of the alloys from 3 parts of alloy 1 and 1 part of alloy 2,
that means that 3/4 of the mixture is made from alloy 1 and 1/4 is made from alloy 2.
Thus the fraction of copper will be 3/4*1/5 + 1/4*4/5 = 7/20 = 35 %
and the fraction of tin will be 1/4*1/5 = 1/20 = 5 %